OpenStudy (anonymous):

Tony is in charge of the concession stand. He charges $4.00 for two hotdogs and two sodas. He charges $4.50 for three hotdogs and one soda. How much does Tony charge per hotdog? How much per soda? I already have my answer, I just want to check it! :)

OpenStudy (stamp):

We are going to use the variables h for hotdogs, and s for soda. $4 is equal to 2 hotdogs and 2 sodas. OR 4 = 2h + 2s also, $4.50 is equal to 3 hotdogs and 1 soda. OR 4.50 = 3h + s -----------------------

OpenStudy (stamp):

You are studying systems of equations. Use either substitution or elimination method to solve for either variable, h or s. Once you have solved for one of the variables, you can use your first variable solution to solve the second variable.

OpenStudy (anonymous):

I got $1.75 per hotdog and $0.75 per soda. Is that correct?

OpenStudy (stamp):

I did not solve the problem, but I can help you check your answer without having ever solved it myself. You say a hotdog h =1.75 and a soda s = 0.75 Our two equations produced were: 4 = 2h + 2s AND 4.5 = 3h + s Does 4 = 2(1.75) + 2(0.75)? Does 4.5 = 3(1.75) + (0.75)?